OpenStudy (anonymous):

help?? Find the first six terms of the sequence. a1 = -8, an = 5 • an-1

OpenStudy (anonymous):

\[ a_2 = 5\cdot a_{2-1} = 5\cdot a_1 = 5(-8) = -40 \]

OpenStudy (anonymous):

\[ a_3 = 5\cdot a_2 = 5(-40) = -200 \]

OpenStudy (anonymous):

Can you finish it?

OpenStudy (anonymous):

They want you to find \(a_1,a_2,a_3,a_4,a_5,a_6\).

OpenStudy (anonymous):

so I keep repeating the formula? the next one would be -1000 correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thanks so much